OpenStudy (anonymous):

Elbert purchased a living room set for $5,021 using a 12-month deferred payment plan. The interest rate after the introductory period is 15.60%. A down payment of $375 is required as well as a minimum monthly payment of $122. What is the balance after the introductory period if only the minimum payment is made until then?

OpenStudy (anonymous):

375 + (12)122 = 375 + 264 = 639 5021 - 639 = 4382 You add the down payment plus the minimum monthly payments of 122 for 12 months, which equals $ 639. You then subtract 639 from $ 5021 to find out how much Elbert has already paid. Therefore, his balance is $ 4382 after the introductory period. He doesn't even have to mess with the 15.60% interest rate until after the introductory period is over, which is in the 13th month. I think I did this correctly, but to make sure you might want to wait for a second opinion.